OpenStudy (anonymous):

simplify the expression. show your work. -3(b-7)

OpenStudy (cggurumanjunath):

-3b+21

OpenStudy (cggurumanjunath):

-3(b-7)=-3*b-3*-7=-3b+21

OpenStudy (anonymous):

thanks man you broke it down well!

OpenStudy (cggurumanjunath):

Anytime. U are always welcome .! :)

OpenStudy (anonymous):

one more if possible... solve the equation. check your answer. I know its something like cross multiplying I think... m/20=5/4

OpenStudy (cggurumanjunath):

100=4m m=100/4 m=25

OpenStudy (anonymous):

ahh pretty simple.. trust me im taking notes.. lol

OpenStudy (cggurumanjunath):

good going !

OpenStudy (anonymous):

one more and ill leave u alone... l x-5 l >3 those l's r the lines on each side..

OpenStudy (anonymous):

can I give u more medals??

OpenStudy (cggurumanjunath):

yes ,u can give more medals if u had posted all 3 questions separately !

OpenStudy (anonymous):

ok ill know next time.. ill look 4 u!!

OpenStudy (cggurumanjunath):

| x- 5 | >3 (x-5) >3 and -(x-5)>3 (x>(3+5)) and (5-x) >3 x>8 and +x +x 5 >(x+3) 5-3>x 2>x

OpenStudy (cggurumanjunath):

have u fanned me ?

OpenStudy (anonymous):

no how do I do that

OpenStudy (cggurumanjunath):

u got it. :)

OpenStudy (anonymous):

i just fanned u!! thanks again really! making progress to get my high school diploma!!

OpenStudy (cggurumanjunath):

ok.

OpenStudy (cggurumanjunath):

good luck buddy .

OpenStudy (anonymous):

thanks!! c ya nxt time!

OpenStudy (cggurumanjunath):

:)